OpenStudy (anonymous):

working together, Paul and Daniel can pick forty bushels of apples in 4.95 hours. Had he done it alone it would have taken Daniel 9 hours. Find how long it would take Paul to do it alone.

hartnn (hartnn):

if a person can do work in x hrs, and another can do it in y hours then they both can do it together in \(\dfrac{1}{1/x+1/y}\) hours.. which is 4.95 here.....

hartnn (hartnn):

wanna know how to get that ?

hartnn (hartnn):

daniel can do the work in 9 hours, so in unit time, (1 hour) how much of total work can he do ?

OpenStudy (anonymous):

u can use this : 1/t1+1/t2+1/t3+...=time work together :) Now we have 4/95 = work together and 1/9 = 1/t1 OK ?

hartnn (hartnn):

how u got 4 ?

OpenStudy (anonymous):

@Missy94

hartnn (hartnn):

what you can do is , plug in x= 9, and find y in (1/x+ 1/y) = 1=4.95

hartnn (hartnn):

***(1/x+ 1/y) = 1/4.95

OpenStudy (anonymous):

amm.. I had another mean:)

OpenStudy (anonymous):

@Missy94 :Do u need help ?

hartnn (hartnn):

daniel can do the work in 9 hours, so in unit time, (1 hour) how much of total work can he do ?

OpenStudy (anonymous):

an hour?

hartnn (hartnn):

yeah, in one hour he will do 1/9 th of the work, right ?

OpenStudy (anonymous):

ya

hartnn (hartnn):

let pauld do the work in x hours, so by same logic how much of work will he do in one hour ?

OpenStudy (anonymous):

1/9?

hartnn (hartnn):

its 1/x total work in x hours, so unit work in 1/x hours together they will do 1/9 +1/x of the work in one hour

hartnn (hartnn):

but in one hour 1/4.95 of the work is done so, 1/9+1/x = 1/4.95 you just need to solve this for x

OpenStudy (anonymous):

11 hours?

hartnn (hartnn):

i got 11 too :)

hartnn (hartnn):

correct.